OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (anonymous):

A number is the solution when the equation is true, i.e. the equality is satisfied (the left and right hand sides are equal) with that number. Example, for the simple equation 2x - 1 = 0, the solutions are x=1/2. To verify this, calculate the left side, it equals 2*(1/2)-1=0, which is equal to the right side, 0. Therefore, x=1/2 is a solution to the equation.